Water leaves a hose at a rate of 1.5kg/s with a speed of 20m/s and is aimed at the side of a car which stops it.Assuming there is no splashing back. What is the force exerted by the water on the car
1
Expert's answer
2021-03-12T07:14:02-0500
"F\Delta t=m\Delta v,""F\Delta t=mv_f-mv_i=-mv_i,""F=\dfrac{-mv_i}{\Delta t}=\dfrac{-1.5\ kg\cdot20\ \dfrac{m}{s}}{1\ s}=-30\ N."

The sign minus means that the force exerted by the water on the car directed in the opposite direction (away from the car).